OpenStudy (anonymous):

If y=sqrt(10-x^2) + sin^-1(12x) then what is the derivative of y. Medal will be given for help, thank you!

OpenStudy (bibby):

split it up \( y=\sqrt{(10-x^2)} + sin^{-1}(12x)\) \( \frac{dy}{dx}=\frac{d}{dx}\sqrt{(10-x^2)} + \frac{d}{dx}sin^{-1}(12x)\)

OpenStudy (anonymous):

Hmm okay, I can handle doing d/dx for sin^1(12x) but how would I find the derivative for sqrt(10-x^2)?

OpenStudy (bibby):

chain rule. sqrt(u) * du

OpenStudy (anonymous):

ahh okay, thank you. I will post my answer shortly if you'd like to confirm it with me.

OpenStudy (bibby):

you can also use wolfram if you want to double check me http://www.wolframalpha.com/input/?i=derive+sqrt%2810-x%5E2%29+%2B+arcsin%2812x%29+

OpenStudy (anonymous):

Sure, thank you

OpenStudy (bibby):

I'll be away for a few minutes so I'll write out the first bit.